If 4 coins are tossed what is the probability of getting 3 heads?
C1, C2, C3, C4
The number of combinations is 4 base 2, or 16.
There are 4 possible scenarios where you get 3 heads.
So the odds are 4:16, or 1:4.
If five dice are rolled how many possible outcomes?
Depends on what detail of outcome you are interested in. If you are only interested in the total, you can get all numbers between 5 (5 ones) and 30 (5 sixes) - a total of 26 different results.
On a dice which number is opposite 3?
Singular of "dice" is "die". Dice are usually constructed so that opposite numbers add up to 7.
Sam had 40 marbles he lost 15 percent of them how many does he have left?
figure it out yourself or ask your mom
In an ordinary 52 card deck, there are 4 three's, and 13 clubs, one of which is the 3 of clubs, so there are 16 cards that are a three or a club. The probability of drawing a three or a club is 16 out of 52, or 4 out or 13, or 0.308.
0
No blue marbles in the bag.
When finding the probability of two or more events what should you do with the fractions?
If it's like this:
"There are 4 books in a bag and two of the four are fiction. Jess randomly pulls one out of the bag and replaces it. What is the probability of getting two fiction books?"
Then you should do this, 2/4 x 2/4 = 4/16 or 1/4
If it's like this:
"There are 4 books in a bag and two of the four are fiction. Jess randomly pulls one out of the bag and doesn't replace it. What is the probability of getting two fiction books?"
Then you should do this, 2/4 x 1/3 = 2/12 or 1/6
What are the outcomes if a die is rolled 100 times?
If you roll a fair, six sided die 100 times, you would expect to see each face one out of six times, or about 17 times.
Statistically, however, you will see variation in the results, because 100 trials is far, far too small to even come close to the theoretical distribution.
If you have a variable X that is normally distributed with mean m and variance s2 then the z-score, Z = (X - m)/s.
Z has a standard Normal distribution.
110
The probability that it is a spade when drawn is 13/52 or 1/4.
The probability of drawing an Ace from a 52 card deck is 4/52 or 1/13.
What is the probability of drawing a queen?
The probability of drawing a queen, from a 52 card deck, is 4/52 or 1/13.
What is probability if odds are 5 to 3?
For 5 to 3 odds of winning; the probability of winning is 0.375 or 37.5%
How many elements are in the event space if you roll a sum of 8 with 2 dice?
There are 5 elements in the event space if you roll a sum of 8 with 2 dice; 2,6 & 6,2 & 3,4 & 4,3 & 4,4.
What are the odds against drawing a queen of spades?
Well in a normal deck of card theres 56, and theres 4 diffrent queens so about, one out of 56. Hope that helps! :)
Here P and S are fixed, so 10 letters are left, out of which 2 are T's
So, 10 letters out of which 2 are T's can be arranged in 10!/2! = 1814400 ways
Now letters P and S can be arranged so that there are 4 letters between them, which can be done in 2*7 = 14 ways
Required no. of ways = 1814400 * 14 = 25401600
6
A deck of cards is shuffled. What is the probability that all spades are together?
This would be like drawing a spade (or any suit for that matter) from a deck of cards one-by-one. Probability is: 13/52*12/51*11/50*10/49*9/48*8/47*7/46*6/45*5/44*4/43*3/42*2/41*1/40 = 6227020800/3.95424E21= 1.57477E-12= 0+
There are 62 or 36 outcomes for rolling 2 dice. The outcomes that meet the criteria specified are: 1,6 2,6 3,6 4,6 6,4 6,3 6,2 6,1. So the probability of one die shows a 6 and the other is a number less than 5 is 8/36 or 2/9.
What is the probability of what?
Guessing them all correctly?
Getting half of the correct?
Getting them all wrong?
PLEASE be specific with your questions if you want WikiAnswers to help.